# Spatial correlations in SIS processes on regular random graphs

### v1

@@ V Computing Shell Probabilities on Regular Random Graphs
Alongside the dynamical trajectory of infection density, we now compute the steady-state value of infection density, $\bar{x}_{I}$, within the corrected mean-field framework, which corresponds to the so-called endemic state. This is of particular interest in percolation theory, where the non-equilibrium phase transition between endemicity and elimination of the virus at a critical rate of infection, $\beta_{c}$, is not fully understood on a wide range of network topologies. Setting $\mathrm{d}x_{I}/\mathrm{d}t=0$ in Eq. (6), we find
@@
where $\bar{x}_{I}^{\text{MF}}=1-\gamma/(\beta k_{0})\equiv 1-1/R_{0}$ is the standard mean-field prediction for the endemic state, obtained by finding the steady state of Eq. (6) with $F_{II}^{(1)}(t)=1$, while $\bar{F}_{II}^{(1)}$ is the steady-state value of the distance-one pairwise correlation function.
For a RRG, $\bar{F}_{II}^{(1)}$ is obtained by setting $\mathrm{d}F_{\ell}\left(t\right)/\mathrm{d}t=0$ in Eq. (8) and using the the expression for $\bar{x}_{I}^{\text{MF}}$, yielding
$\bar{F}_{II}^{(\ell)}=\frac{\bar{F}_{II}^{(1)}-\bar{F}_{II}^{(\ell)}\bar{x}_{I}^{\text{MF}}}{\bar{F}_{II}^{(1)}-\bar{x}_{I}^{\text{MF}}}\!\!\!\!\!\!\sum_{\ell^{\prime}=\max\{\ell-1,1\}}^{\ell+1}\!\!\!\!\!\!\!\!\!\!p\left(\ell^{\prime}|\ell\right)\bar{F}_{II}^{(\ell^{\prime})}+\frac{\bar{F}_{II}^{(1)}\delta_{\ell,1}}{k_{0}\bar{x}_{I}^{\text{MF}}}.$ (15)
@@
where $\alpha$ and $K=\exp(-\lambda)$ are parameters yet to be determined. Employing this empirical dependence allows finding $\bar{F}_{II}^{(\ell)}$. Substituting Eq. (16) into (15) for shells $\ell=1,2$ and further using Eq. (16) for $\ell=3$, yields two implicit equations for $\alpha$ and $K$. Following algebraic manipulations detailed in Appendix A, these equations reduce to a cubic equation $AK^{3}+BK^{2}+CK+D=0$, alongside an explicit equation connecting $\alpha$ and $K$, which can be solved numerically for any $k_{0}$. Notably, while Eq. (15) breaks down at larger values of $\ell$ (see Fig. 5(b)), it remains rather accurate for $\ell={\cal O}(1)$; this allows for an accurate approximation of the distance-one pairwise correlation, which is the primary quantity of interest in our analysis, as it determines the correction to $\bar{x}_{I}$.
@@
Figure 5: (a)-(b) The steady-state pairwise correlation function, $\bar{F}_{II}^{\ell}$, versus the network distance $\ell$, on linear and semi-logarithmic scales, respectively. Circles indicate values obtained from numerical simulations, while solid lines represent exponential fits based on Eq. (16). (c)-(d) The steady-state pairwise correlation function $\bar{F}_{II}^{1}$ and steady-state infected fraction $\bar{x}_{I}$, respectively, versus the average node degree $k_{0}$. Numerical solutions (circles) and simulation values (triangles) show strong agreement. Additionally, power series approximations of the second order of $k_{0}^{-1}$ (solid green lines) converge toward the numerical and simulation results as $k_{0}$ increases.
@@
The analysis above greatly simplifies under the assumption of $k_{0}\ll N^{1/2}$, making the network locally tree like  [^1] In the complementary regime $k_{0}\gtrsim N^{1/2}$, the regular random graph is no longer locally tree-like and finite-size effects as well as an increased prevalence of short cycles in the network occurs. However, since the expected number of shells in a RRG scales as $\log_{k_{0}}\!N$, for $k_{0}\gtrsim N^{1/2}$ the network has only $\sim 2$ shells, which is clearly a less interesting regime to deal with.. In this limit, the conditional probabilities read
$p(0|0)$ $=0,\quad p(1|0)=1,$
$p(0|1)$ $=1/k_{0},\quad p(1|1)=\mathcal{O}(k_{0}/N),$
$p(2|1)$ $=1-1/k_{0}+\mathcal{O}(k_{0}/N),$
$p(1|2)$ $=1/k_{0}+\mathcal{O}(k_{0}/N),\quad p(2|2)=\mathcal{O}(k_{0}^{2}/N),$
$p(3|2)$ $=1-1/k_{0}+\mathcal{O}(k_{0}^{2}/N).$ (17)
Substituting these results into Eq. (15) together with the ansatz from Eq. (16), we obtain
$\alpha=-\frac{1}{\left(k_{0}-1\right)\left(R_{0}-1\right)K\left(K-R_{0}\right)+R_{0}K},$ (18)
$0=K^{3}-R_{0}\frac{\left(2k_{0}-1\right)}{k_{0}}K^{2}+\left[R_{0}^{2}-\frac{R_{0}-1}{k_{0}-1}\right]K-\frac{R_{0}}{k_{0}}.$ (19)
Further assuming $k_{0}\gg 1$, and expanding $K$ and $\alpha$ in powers of $k_{0}^{-1}$ in the regime $1\ll k_{0}\ll N^{1/2}$ gives
$K$ $=\frac{1}{R_{0}k_{0}}+\frac{R_{0}+1}{R_{0}^{3}k_{0}^{2}}+\mathcal{O}\left(\frac{1}{k_{0}^{3}}\right),$ (20)
$\alpha$ $=\frac{1}{R_{0}-1}+\frac{R_{0}^{2}-R_{0}+1}{R_{0}(R_{0}-1)^{2}k_{0}}+\mathcal{O}\left(\frac{1}{k_{0}^{2}}\right).$ (21)
@@
Here, we have retained terms in the series expansion of $\alpha$ up to ${\cal O}(k_{0}^{-1})$, since the leading term in the expansion of $K$ is also of ${\cal O}(k_{0}^{-1})$, while we are interested in expressing their product $\bar{F}_{II}^{(1)}=1+\alpha K$ up to ${\cal O}(k_{0}^{-2})$. This yields
$\bar{F}_{II}^{(1)}\!=\!1\!+\!\frac{1}{(R_{0}\!-\!1)R_{0}k_{0}}\!+\!\frac{R_{0}^{3}+R_{0}-1}{(R_{0}\!-\!1)^{2}R_{0}^{3}k_{0}^{2}}\!+\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\!.$ (22)
Using (14), this yields the steady-state infection density
$\bar{x}_{I}=\bar{x}_{I}^{\text{MF}}\left[1\!-\!\frac{1}{R_{0}\left(R_{0}\!-\!1\right)k_{0}}\!-\!\frac{R_{0}^{3}-1}{R_{0}^{3}\left(R_{0}\!-\!1\right)^{2}\!k_{0}^{2}}\!-\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\!\right]\!\!.$ (23)
@@
Comparisons between our predictions for $\bar{F}_{II}^{(1)}$ and $\bar{x}_{I}$ and numerical simulations are shown in Figs. 5(c)-(d) and excellent agreement is observed at large $k_{0}$. Notably, these expansions make explicit how finite-degree corrections alter both the two-point correlation function and infection density relative to the standard mean-field predictions. The consistent negativity of the correction terms in $\bar{x}_{I}$ reflects the dampening effect of local clustering and correlations on disease spread in finite-connectivity networks, consistent with the fact that $\bar{F}_{II}^{(1)}(t)\geq 1$ at all distances, as seen in Fig. 2.
@@
Finally, exploiting the power series expansions of the steady-state infection density $\bar{x}_{I}$ and pairwise correlation function $\bar{F}_{II}^{(1)}$, we also performed a linear stability analysis of the coupled epidemic dynamics given by Eqs. (6) and (8) for $\ell=1$, assuming that the relaxation time of all $\bar{F}_{II}^{(\ell)}$ is comparable to that of $\bar{F}_{II}^{(1)}$, for $\ell>1$. By evaluating the Jacobian matrix at the endemic equilibrium, we explicitly capture corrections beyond the mean-field limit. Expanding the resulting relaxation time (corresponding to the lowest eigenvalue) up to ${\cal O}(k_{0}^{-2})$ we find
$\tau=\frac{1}{R_{0}\!-\!1}\!\!\left[1\!+\!\frac{4R_{0}-2}{\left(R_{0}\!-\!1\right)^{2}k_{0}}\!+\!\frac{2\left(4R_{0}\!-\!7\right)R_{0}^{3}\!+\!2}{\left(R_{0}\!-\!1\right)^{3}R_{0}^{2}k_{0}^{2}}\!+\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\right]\!.$ (24)
The zeroth-order value $1/(R_{0}-1)$ is the mean-field relaxation time, while the higher-order positive corrections imply that for lower $k_{0}$, the system takes more time to converge.
@@ VII Discussion
In this work, we have characterized, through numerical simulations and analytical calculations, the role of spatial correlations in the transmission dynamics of SIS processes across various network topologies. Our findings demonstrate that spatial correlations in infectiousness are suppressed by decreasing the average shortest path length by random edge rewiring and increasing the infection rate. The observed sensitivity of spatial correlations to network topology implies that corrections to mean-field theory previously derived for regular lattices are unsuitable for RRGs. To address this, we leveraged existing analytical results on RRG connectivity to develop appropriate mean-field corrections tailored specifically to regular random topologies. These new corrections provide significantly improved predictions of infection dynamics in SIS processes on RRGs, particularly for values of the basic reproduction number above the epidemic threshold (i.e., $R_{0}>1$). The transition at $R_{0}=1$ defines the critical epidemic threshold separating disease extinction from persistence. Traditionally, the mean-field approximation predicts a steady-state infected fraction $\bar{x}_{I}^{\text{MF}}=1-1/R_{0}$. However, our corrections adjust this estimate and reflect spatial correlations arising from clustering of infections. This adjustment reduces the predicted endemic prevalence, capturing the clustering effect of infected nodes. As a result, the effective pool of susceptible neighbors per infected individual is smaller than assumed in mean-field theory, which in turn lowers the steady-state infection density.
@@
While the corrected mean-field approach outperforms previous lattice-based corrections in this regime, the accuracy of the predictions still depends on the proximity of the rate of infection, $\beta$, to the critical value for a given underlying network structure. Notably, the accuracy of these corrections diminishes near the critical threshold, indicating the need for more robust approximation methods in near-critical regimes.
@@
Future research could attempt to re-derive corrections for the RRG by invoking the Fisher-Kopeliovich approximation, which is known to provide improved accuracy where the KSA becomes inadequate [21] (albeit at the expense of greater analytical complexity). Another potential direction would be to expand the current framework to topologies on which node degree is heterogeneous. This will enable us to predict and analyze more realistic scenarios, particularly where degree variability plays a crucial role in epidemic outcomes [22].
@@
Overall, our work broadens the applicability of existing corrections to mean-field theory and highlights how network structure impacts spatial correlations in SIS processes. A promising direction for future research involves exploring intermediate network structures to bridge the analytical gap between regular lattices and fully randomized network topologies.
Appendix A Corrections to the steady state density
@@
Here we derive closed-form equations for $\alpha$ and $K$ in the main text, which allow finding the correction to the steady-state infection density for arbitrary $k_{0}$. Writing Eq. (15) for shells $\ell=1,2$ yields
@@
and
$0=p\left(3|2\right)\bar{x}_{I}^{\text{MF}}K^{3}+\left(\bar{x}_{I}^{\text{MF}}p\left(2|2\right)-p\left(3|2\right)\right)K^{2}$ (26)
@@
Finally, substituting the value of $\alpha$ from Eq. (25) into Eq. (26) and rearranging terms yields a cubic equation for $K$, $AK^{3}+BK^{2}+CK+D=0$, where the coefficients satisfy:
$A$ $=k_{0}{x}_{I}^{\text{MF}2}p\left(3|2\right)\left(1-p\left(1|1\right)\right)$
$\hphantom{=}-k_{0}{x}_{I}^{\text{MF}}p\left(3|2\right)p\left(2|1\right)-p\left(3|2\right),$ (27)
$B$ $=k_{0}{x}_{I}^{\text{MF}2}\left[p\left(2|2\right)-p\left(1|1\right)p\left(2|2\right)+p\left(1|1\right)p\left(3|2\right)\right]$
$\hphantom{=}+k_{0}{x}_{I}^{\text{MF}}p\left(2|1\right)\left[1+p\left(3|2\right)-p\left(2|2\right)\right]-p\left(2|2\right){x}_{I}^{\text{MF}},$ (28)
$C$ $=k_{0}{x}_{I}^{\text{MF}2}\left[p\left(1|2\right)-p\left(1|1\right)p\left(1|2\right)+p\left(1|1\right)p\left(2|2\right)\right]$
$\hphantom{=}+k_{0}{x}_{I}^{\text{MF}}\left[1-p\left(2|1\right)\left(1-p\left(2|2\right)+p\left(1|2\right)\right))\right]$
$\hphantom{=}-p\left(1|2\right){x}_{I}^{\text{MF}}+\frac{{x}_{I}^{\text{MF}}\left(1-k_{0}\right)}{1-{x}_{I}^{\text{MF}}},$ (29)
$D$ $=k_{0}{x}_{I}^{\text{MF}2}p\left(1|1\right)p\left(1|2\right)+k_{0}{x}_{I}^{\text{MF}}p\left(2|1\right)p\left(1|2\right).$ (30)

### v2

@@ VI A modified steady-state analysis on random regular networks
Alongside the dynamical trajectory of infection density, we now compute the steady-state value of infection density, $\bar{x}_{I}$, predicted within the MPM framework. This quantity is of particular interest in percolation theory, where the non-equilibrium phase transition between endemicity and disease elimination at a critical infection rate, $\beta_{c}$, is not fully understood on a wide range of network topologies. Setting $\mathrm{d}x_{I}/\mathrm{d}t=0$ in Eq. (6), we find
@@
where $\bar{x}_{I}^{\text{MF}}=1-\gamma/(\beta k_{0})\equiv 1-1/\Lambda$ is the standard mean-field prediction for the endemic state, obtained by finding the steady-state value of Eq. (6) with $F_{II}^{(1)}(t)=1$, while $\bar{F}_{II}^{(1)}$ is the steady-state value of the distance-one pairwise correlation function. The ratio $\Lambda=\beta k_{0}/\gamma$ is known as the effective spreading rate [16, 15], a dimensionless control parameter that dictates the epidemic threshold $\Lambda=1$ in the well-mixed limit.
For an RRG, $\bar{F}_{II}^{(\ell)}$ is obtained by setting $\mathrm{d}F_{II}^{(\ell)}/\mathrm{d}t=0$ in Eq. (8) and using the expression for $\bar{x}_{I}^{\text{MF}}$, yielding
$\hskip-8.5359pt\bar{F}_{II}^{(\ell)}\!=\!\frac{\bar{F}_{II}^{(1)}\!-\!\bar{F}_{II}^{(\ell)}\bar{x}_{I}^{\text{MF}}}{\bar{F}_{II}^{(1)}-\bar{x}_{I}^{\text{MF}}}\!\!\!\!\!\!\!\sum_{\ell^{\prime}=\max\{\ell-1,1\}}^{\ell+1}\!\!\!\!\!\!\!\!\!\!\!\!p\left(\ell^{\prime}|\ell\right)\bar{F}_{II}^{(\ell^{\prime})}+\frac{\bar{F}_{II}^{(1)}\delta_{\ell,1}}{k_{0}\bar{x}_{I}^{\text{MF}}}.$ (15)
@@
where $\alpha$ and $K=\exp(-\lambda)$ are parameters yet to be determined. This empirical ansatz allows one to approximate $\bar{F}_{II}^{(\ell)}$. Substituting Eq. (16) into (15) for shells $\ell=1,2$ and further using Eq. (16) for $\ell=3$, yields two implicit equations for $\alpha$ and $K$. Following algebraic manipulations detailed in Appendix B, these equations reduce to a cubic equation $AK^{3}+BK^{2}+CK+D=0$, alongside an explicit equation connecting $\alpha$ and $K$, which can be solved numerically for any $k_{0}$. Notably, while Eq. (15) breaks down at larger values of $\ell$ (see Fig. 5(b)), it remains rather accurate for $\ell={\cal O}(1)$; this allows for an accurate approximation of the distance-one pairwise correlation, which is the primary quantity of interest in our analysis, as it determines the correction to $\bar{x}_{I}$.
@@
Figure 5: (a)-(b) The steady-state pairwise correlation function, $\bar{F}_{II}^{(\ell)}$, versus the network distance $\ell$, on linear and semi-logarithmic scales, respectively, for $\Lambda=1.7$. Circles indicate values obtained from numerical simulations, while solid lines are exponential fits based on Eq. (16). (c)-(d) $\bar{F}_{II}^{(1)}$ and $\bar{x}_{I}$, respectively, versus the average degree $k_{0}$: comparison between numerical solutions (circles), simulation results (triangles), and the large-$k_{0}$ asymptotes (solid green lines), see text.
@@
The analysis above greatly simplifies under the assumption of $k_{0}\ll N^{1/2}$, making the network locally tree-like  [^1] In the complementary regime $k_{0}\gtrsim N^{1/2}$, the RRG is no longer locally tree-like and finite-size effects as well as an increased prevalence of short cycles in the network occurs. Yet, as the expected number of shells scales as $\log_{k_{0}}\!N$, for $k_{0}\gtrsim N^{1/2}$ the network has only $\sim 2$ shells, which is clearly a less interesting regime to deal with.. In this limit, the conditional probabilities read
$p(0|0)\!=\!0,\;\;p(1|0)\!=\!1,\;\;p(0|1)\!=\!1/k_{0},\;\;p(1|1)=\mathcal{O}(k_{0}/N),$
$p(2|1)\!=\!1-1/k_{0}+\mathcal{O}(k_{0}/N),\;\;p(1|2)\!=\!1/k_{0}+\mathcal{O}(k_{0}/N),$
$\hskip-34.1433ptp(2|2)\!=\!\mathcal{O}(k_{0}^{2}/N),\;\;p(3|2)\!=\!1-1/k_{0}+\mathcal{O}(k_{0}^{2}/N).$ (17)
Substituting these results and Eq. (16) into (15), yields
$\alpha=-\left[\left(k_{0}-1\right)\left(\Lambda-1\right)K\left(K-\Lambda\right)+\Lambda K\right]^{-1},$
$0=K^{3}-\Lambda\frac{\left(2k_{0}-1\right)}{k_{0}}K^{2}+\left[\Lambda^{2}-\frac{\Lambda-1}{k_{0}-1}\right]K-\frac{\Lambda}{k_{0}}.$ (18)
@@
Further assuming $k_{0}\gg 1$, and expanding $K$ and $\alpha$ in powers of $k_{0}^{-1}$ in the regime $1\ll k_{0}\ll N^{1/2}$ gives $K=1/(\Lambda k_{0})+(\Lambda+1)/(\Lambda^{3}k_{0}^{2})+\mathcal{O}\left(k_{0}^{-3}\right)$, and $\alpha=1/(\Lambda-1)+(\Lambda^{2}-\Lambda+1)/[\Lambda(\Lambda-1)^{2}k_{0}]+\mathcal{O}\left(k_{0}^{-2}\right)$. Here, we have retained terms in the series expansion of $\alpha$ up to ${\cal O}(k_{0}^{-1})$, since the leading term in the expansion of $K$ is also of ${\cal O}(k_{0}^{-1})$, while we are interested in expressing their product $\bar{F}_{II}^{(1)}=1+\alpha K$ up to ${\cal O}(k_{0}^{-2})$. This yields
$\bar{F}_{II}^{(1)}\!=\!1\!+\!\frac{1}{(\Lambda\!-\!1)\Lambda k_{0}}\!+\!\frac{\Lambda^{3}+\Lambda-1}{(\Lambda\!-\!1)^{2}\Lambda^{3}k_{0}^{2}}\!+\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\!.$ (19)
Therefore, using Eq. (14), we obtain
$\bar{x}_{I}=\bar{x}_{I}^{\text{MF}}-\frac{1}{k_{0}\Lambda^{2}}-\frac{\Lambda^{2}+\Lambda+1}{k_{0}^{2}\Lambda^{4}}-\mathcal{O}\left(\frac{1}{k_{0}^{3}}\right).$ (20)
@@
Our predictions for $\bar{F}_{II}^{(1)}$ and $\bar{x}_{I}$ agree well with numerical simulations, see Figs. 5(c)-(d). These expansions make explicit how finite-degree corrections alter both the two-point correlation function and infection density relative to the standard mean-field predictions. The fact that the correction terms in $\bar{x}_{I}$ are negative reflects the dampening effect of local clustering and correlations on disease spread in finite-connectivity networks, consistent with $\bar{F}_{II}^{(1)}(t)\geq 1$ at all distances, as seen in Fig. 2. Notably, equating Eq. (20) to zero yields the critical spreading rate
$\Lambda_{c}=1+k_{0}^{-1}+2k_{0}^{-2}+\mathcal{O}\left(k_{0}^{-3}\right),$ (21)
which differs from the SPM threshold at order $\mathcal{O}(k_{0}^{-2})$, see Appendix C for a detailed comparison between the MPM and SPM  [^2] Note that, once the corrected critical spreading rate $\Lambda_{c}$ is computed, the basic reproduction number can be readily obtained as $R_{0}=\Lambda/\Lambda_{c}$..
@@
Finally, exploiting the power series expansions of the steady-state infection density $\bar{x}_{I}$ and pairwise correlation function $\bar{F}_{II}^{(1)}$, we also perform a linear stability analysis of the coupled epidemic dynamics given by Eqs. (6) and (8) for $\ell=1$, assuming that the relaxation time of all $\bar{F}_{II}^{(\ell)}$, for $\ell>1$, is comparable to that of $\bar{F}_{II}^{(1)}$. By evaluating the Jacobian matrix at the endemic equilibrium, we explicitly capture corrections beyond the mean-field limit. Expanding the resulting relaxation time (corresponding to the eigenvalue with the smallest magnitude) up to ${\cal O}(k_{0}^{-1})$ we find $\tau_{\mathrm{relax}}\simeq 1/(\Lambda-1)\left[1+2(2\Lambda-1)/(\left(\Lambda-1\right)^{2}k_{0})\right]$. The zeroth-order value $1/(\Lambda-1)$ is the mean-field relaxation time, while higher-order positive corrections imply that for lower $k_{0}$, the system takes more time to converge.
@@ VII Discussion
In this work, we have characterized, through numerical simulations and analytical calculations, the role of spatial correlations in the transmission dynamics of SIS processes across various network topologies. Our findings demonstrate that spatial correlations between infectious nodes are suppressed both by random edge rewiring, which decreases the average shortest-path length, and by increasing the infection rate. The observed sensitivity of spatial correlations to network topology implies that corrections to mean-field theory previously derived for regular lattices are unsuitable for RRGs. To address this, we leveraged existing analytical results on RRG connectivity to develop appropriate mean-field corrections tailored specifically to random regular topologies. These corrections provide improved predictions of infection dynamics of SIS processes on RRGs, when the effective spreading rate is above the epidemic threshold, i.e., $\Lambda>\Lambda_{c}$. Traditionally, the well-mixed approximation assumes this critical transition occurs at $\Lambda=1$, predicting a steady-state infected fraction $\bar{x}_{I}^{\text{MF}}=1-1/\Lambda$. Our corrections refine this estimate by explicitly accounting for the spatial clustering of infected nodes. Thus, the effective pool of susceptible neighbors per infected individual is smaller than that assumed in mean-field theory, thereby lowering the predicted $\bar{x}_{I}^{\text{MF}}$. Our results are also superior compared to the SPM, especially at low average degrees where multi-shell correlations are most significant (see Appendix C).
@@
While the MPM outperforms previous lattice-based corrections in this regime, the accuracy of the predictions still depends on the proximity of the rate of infection, $\beta$, to the critical value for a given underlying network structure. Notably, the accuracy of these corrections diminishes near the critical threshold, indicating the need for more robust approximation methods close to criticality.
@@
Future research could attempt to re-derive corrections for the RRG by invoking the Fisher-Kopeliovich approximation, which is known to provide improved accuracy where the Kirkwood superposition approximation becomes inadequate [14] (albeit at the expense of greater analytical complexity). Another potential direction would be to expand the current framework to heterogeneous topologies. This will enable us to predict and analyze more realistic scenarios, particularly where degree variability plays a crucial role in epidemic outcomes [6].
@@
Overall, this work broadens the applicability of existing corrections to mean-field theory and highlights how network structure impacts spatial correlations in SIS processes. Another promising future direction involves exploring network structures with intermediate randomness to bridge the analytical gap between regular lattices and random regular topologies.
Appendix A Edge-swapping algorithm
Here, we present an edge swapping algorithm which is a modification of an algorithm from Wang et al. [20]. Given a network $G(V,E)$ representing a regular lattice, where $E$ and $V$ are the edge and node sets, the algorithm can be summarized as follows:
  • If $p_{e}\leq p$, swap the edges $e_{2i-1}=(u_{2i-1},v_{2i-1})$ and $e_{2i}=(u_{2i},v_{2i})$ to $e_{2i-1}^{\prime}=(u_{2i-1},v_{2i})$ and $e_{2i}^{\prime}=(u_{2i},v_{2i-1})$; otherwise, keep the edges as they are, $e_{2i-1}^{\prime}=e_{2i-1}$ and $e_{2i}^{\prime}=e_{2i}$.
This yields a revised network, $G=(V,E^{\prime})$, with $E^{\prime}\!=\!\{e_{1}^{\prime},e_{2}^{\prime},\dots,e_{|E|}^{\prime}\}$. Notably, these randomized network structures converge to the RRG as $p$ becomes ${\cal O}(1)$.
Appendix B Corrections to the steady-state density
@@
Here we derive closed-form equations for $\alpha$ and $K$ in the main text, which allow one to determine the correction to the steady-state infection density for arbitrary $k_{0}$. Writing Eq. (15) for shells $\ell=1,2$ yields
@@
$0=p\left(3|2\right)\bar{x}_{I}^{\text{MF}}K^{3}+\left(\bar{x}_{I}^{\text{MF}}p\left(2|2\right)-p\left(3|2\right)\right)K^{2}$
@@
Substituting $\alpha$ from the first of Eqs. (22), yields a cubic equation for $K$, $AK^{3}\!+\!BK^{2}\!+\!CK\!+\!D\!=\!0$, with
$A$ $=k_{0}(\bar{x}_{I}^{\text{MF}})^{2}p\left(3|2\right)\left(1-p\left(1|1\right)\right)$
$\hphantom{=}-k_{0}\bar{x}_{I}^{\text{MF}}p\left(3|2\right)p\left(2|1\right)-p\left(3|2\right),$
$B$ $=k_{0}(\bar{x}_{I}^{\text{MF}})^{2}\left[p\left(2|2\right)-p\left(1|1\right)p\left(2|2\right)+p\left(1|1\right)p\left(3|2\right)\right]$
$\hphantom{=}+k_{0}\bar{x}_{I}^{\text{MF}}p\left(2|1\right)\left[1+p\left(3|2\right)-p\left(2|2\right)\right]-p\left(2|2\right)\bar{x}_{I}^{\text{MF}},$
$C$ $=k_{0}(\bar{x}_{I}^{\text{MF}})^{2}\left[p\left(1|2\right)-p\left(1|1\right)p\left(1|2\right)+p\left(1|1\right)p\left(2|2\right)\right]$
$\hphantom{=}+k_{0}\bar{x}_{I}^{\text{MF}}\left[1-p\left(2|1\right)\left(1-p\left(2|2\right)+p\left(1|2\right)\right)\right]$
$\hphantom{=}-p\left(1|2\right)\bar{x}_{I}^{\text{MF}}+\bar{x}_{I}^{\text{MF}}\left(1-k_{0}\right)/(1-\bar{x}_{I}^{\text{MF}}),$
$D$ $=k_{0}(\bar{x}_{I}^{\text{MF}})^{2}p\left(1|1\right)p\left(1|2\right)+k_{0}\bar{x}_{I}^{\text{MF}}p\left(2|1\right)p\left(1|2\right).$
Appendix C Comparison with the SPM
The standard pairwise model (SPM) [7, 10] is a well-established and widely accepted closure scheme in network epidemiology, offering a major improvement over mean-field approaches by explicitly accounting for local network structure through nearest-neighbor correlations. The SPM yields the following steady-state value: $\bar{x}_{I}^{\text{SPM}}\!=\!1\!-\!1/\left[\Lambda\!-\!1/(k_{0}\!-\!1)\right]$, where $\Lambda=\beta k_{0}/\gamma$. Expansion of $\bar{x}_{I}^{\text{SPM}}$ in powers of $k_{0}^{-1}\!\ll\!1$ yields
$\bar{x}_{I}^{\text{SPM}}\simeq\bar{x}_{I}^{\text{MF}}-\frac{1}{k_{0}\Lambda^{2}}-\frac{\Lambda^{2}+\Lambda}{k_{0}^{2}\Lambda^{4}},$ (23)
up to $\mathcal{O}(k_{0}^{-3})$ corrections. The $\mathcal{O}(k_{0}^{-1})$ term agrees with our MPM result [Eq. (20)], yet the ${\cal O}(k_{0}^{-2})$ term differs. This reflects finite-degree clustering effects that are not captured by the SPM and is precisely what leads to the smaller errors of the MPM steady-state predictions.
Figure 6(a) presents the steady-state prevalence $\bar{x}_{I}$ across a range of mean degrees $k_{0}$ for $\Lambda=1.7$. The MPM with $\ell_{\max}>1$ (blue) agrees well with simulation data (red triangles), with deviations around $1\!-\!2\%$ over the entire range of $k_{0}$. In contrast, the SPM (dark green line) systematically overestimates the prevalence, with large errors for $k_{0}={\cal O}(1)$. This reflects the contribution of higher-order spatial correlations captured by evaluating the closure at multiple shell distances. Figure 6(b) shows the corresponding steady-state distance-one pairwise correlation function $\bar{F}_{II}^{(1)}$. Here, the MPM again agrees well with simulations, while the SPM greatly underestimates $\bar{F}_{II}^{(1)}$ for $k_{0}={\cal O}(1)$.
Figure 6: (a) Steady-state infection density $\bar{x}_{I}$ and (b) correlation function $\bar{F}_{II}^{(1)}$ versus $k_{0}$ for $\Lambda=1.7$. Shown are the MPM (blue line), SPM (dark green line) and numerical simulations (red triangles) with $N=10^{4}$ nodes. Insets show relative error between simulations and MPM (circles), and SPM (squares).